# LEGO_HQEC: A Software Tool for Analyzing Holographic Quantum Codes

### v1

@@ IV.2 Runtime Performance
As we were limited by computational power, we were able to extract results for only the first two layers using the integer-optimization decoder. For all other decoders and the OperatorPush method, however, we were able to obtain time-scaling estimates for up to $L=4$, i.e. on the order of roughly several thousand boundary (physical) qubits, depending on the Schläfli symbols $\{p,q\}$ associated with the commensurate hyperbolic tiling.
V Conclusion
@@
In this work, we have presented and demonstrated LEGO_HQEC, a new software tool for constructing and testing holographic quantum codes under typical noise channels of interest. Similar to other notable software packages for quantum error correction, fault tolerance, among others [72, 73, 74, 75, 76, 77], our focus with LEGO_HQEC has been to facilitate open-source development and enable future scientific discovery in the realm of holographic quantum error correction. In the process, we have: developed and implemented an easy-to-use open-source software tool; demonstrated several examples of its capabilities; and have provided a basic framework upon which more detailed code studies and further discoveries in holographic quantum error correction can be leveraged.
@@
LEGO_HQEC has already featured in several holographic code studies, including those exploring the novel Evenbly codes [52, 63], analyses of asymptotically zero-rate holographic codes as compared to the hashing bound [55, 65], as well as for constructions of heterogeneous holographic codes [62], which are constructed from several seed codes and can circumvent the Eastin-Knill theorem regarding codes with universal logical gate sets [78, 31, 32].
In [49, 60, 76], tensor-network decoders have been implemented using the Julia language (which is available at https://github.com/qecsim/TensorNetworkCodes.jl). Although one can technically build holographic tensor-network codes using this package, the principal difference lies in the fact that our work is streamlined and specialized for generating regular hyperbolic lattices with a quasiregular boundary in an efficient manner, as well as for testing codes under a wide range of possible noise channels; these include the erasure, depolarizing, and biased-noise channels (such as those explored in [55]).
@@
There are many ideas for future modification of LEGO_HQEC. Firstly, an extension for considering any hyperbolic tiling parameterized by $\{p,q\}$ would be enormously helpful, as it has been shown that certain classes of holographic codes have vastly different threshold behavior as the tiling is altered [52]. Furthermore, recent work [62, 65] shows that incorporating multiple seed codes can allow for potentially universal fault-tolerant logical operations [62]; tools to support such fault-tolerance studies could be very useful. One can also consider implementing functionality for greater control over the encoded logical qubits in the bulk, as recent work shows that recovery rates of finite-rate holographic codes greatly depend not only on the number $k$ of encoded logical qubits, but critically on the positions of said logical qubits in the bulk [79]. Even more broadly, it is known that the qLEGO formalism applies to non-Abelian stabilizer codes, such as the XP stabilizer variety [80, 81], or even non-stabilizer codes, such as those presented in [82, 7, 83, 84, 85]; one may consider incorporating functionality to accommodate such holographic code constructions in the near future. Moreover, the class of codes known as Evenbly codes, a derivate of the hyperinvariant tensor networks first introduced in [86, 87], have been extensively studied using LEGO_HQEC [52, 63]; however, one major impediment was the inclusion of Hadamard gates on the contracted edges of the tensor network, thereby hindering decoding of Pauli channels with the tensor-network decoder. Allowing for such constraints, among others, remains within the future goals of this work. Finally, there are many additions purely from the software and usability standpoint, such as: adding functionality for tree-style concatenated codes [23, 10, 7]; visualization tools; or GPU acceleration tools; among other prospects.
V.1 Acknowledgements
@@
We thank David Elkouss and Rob Harris for useful discussions. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation”. CC acknowledges support from the Commonwealth Cyber Initiative at Virginia Tech.
V.2 Author Contributions

### v3

@@ IV-B Runtime Performance
We were able to extract results for only the first two layers using the integer-optimization decoder, as we were limited by computational power. For all other decoders and the OperatorPush method, however, we were able to obtain time-scaling estimates for up to $L=4$, i.e., on the order of roughly several thousand boundary (physical) qubits, depending on the Schläfli symbols $\{p,q\}$ associated with the commensurate hyperbolic tiling.
V Results for Holographic Black Hole Codes
As a further demonstration of the range of our software tool’s usability, we investigate the threshold behavior of the black hole HaPPY code under the erasure channel, which was introduced in [heterogeneous_holo_qrm].
Figure 8: The black hole HaPPY code with five logical qubits (red legs), constructed by removing the central tensor from the zero-rate code of Figure 3(c). Here, we show the black hole code with (a) $L=2$ layers and (b) $L=3$ layers of edge inflation, corresponding to 95 and 375 physical qubits (outer boundary legs), respectively.
Instances of the code are constructed by removing the central tensor and then growing $L$ edge inflation layers of the hexagon code tiling (see Figure 8). Each instance encodes five logical qubits; stabilizers and logical operators are generated with OperatorPush, and decoding is performed using the Gaussian-elimination (GE) erasure decoder with $10^{5}$ Monte Carlo trials per $(p,L)$. Logical success probabilities are evaluated with respect to the simultaneous recoverability of all five logical qubits. We report a threshold $p_{\text{th}}=50\%$ from finite-size crossings (Figure 9). Consistent with our data, $p_{\text{th}}=50\%$ coincides with that of the zero-rate HaPPY code, while the black hole HaPPY code exhibits uniformly lower success probabilities; moreover, as $L$ increases the rate of the black hole HaPPY code approaches zero. Besides the black hole HaPPY code, we also constructed and tested the black hole holographic Steane code [harris2018calderbank] and the black hole holographic $[[6,1,3]]$ code [farrelly2021tensor], obtaining similar threshold results around $p_{\text{th}}=50\%$. These black hole codes are also asymptotically zero-rate; however, since they encode more logical qubits than their single-logical-qubit counterparts, their recovery curves lie slightly below while maintaining the same threshold.
Figure 9: Recovery probability of the black-hole HaPPY code under the erasure channel as a function of $P_{\text{erasure}}$ for $L\in\{1,\dots,4\}$. Each instance encodes five logical qubits, with success defined as simultaneous recovery of all. The threshold matches with that of the one-logical-qubit zero-rate HaPPY code, but with lower recovery probability. The black hole code is constructed by removing the tensor at layer $L=0$ from a regular holographic code (see Fig. 8); hence, the smallest code starts from $L=1$. For the simulations, $L\in\{1,2,3\}$ used 10,00010{,}000 Monte Carlo trials for each $P_{\text{erasure}}$ value, and $L=4$ used 1,0001{,}000 trials, with step size $\Delta p_{\text{erasure}}=0.01$. The computation times were $15$ s, $141$ s, $989$ s, and $3805$ s for $L\in\{1,\dots,4\}$, respectively (on an i9-14900K using 24 threads).
VI Conclusion
@@
In this work, we have presented and demonstrated LEGO_HQEC, a new software tool for constructing and testing holographic quantum codes under typical noise channels of interest. Similar to other notable software packages for quantum error correction, fault tolerance, among others [qgym_feld, sarkar2024yaqq, wille_qldpc, gidney2021stim, qecsim, QiskitCommunity2017], our focus with LEGO_HQEC has been to facilitate open-source development and enable future scientific discovery in the realm of holographic quantum error correction. In the process, we have: developed and implemented an easy-to-use open-source software tool; demonstrated several examples of its capabilities; and have provided a basic framework upon which more detailed code studies and further discoveries in holographic quantum error correction can be leveraged.
@@
LEGO_HQEC has already featured in several holographic code studies, including those exploring the novel Evenbly codes [steinberg2024far, steinberg_2023], analyses of asymptotically zero-rate holographic codes as compared to the hashing bound [fan2024overcoming, junyu_thesis], as well as for constructions of heterogeneous holographic codes [heterogeneous_holo_qrm], which are constructed from several seed codes and can circumvent the Eastin-Knill theorem regarding codes with universal logical gate sets [eastin2009restrictions, jochym2014using, chamberland2016thresholds].
As a proof of concept, we presented the black-hole HaPPY code and its threshold under the erasure noise channel. Our numerical analysis shows that, while its threshold under the erasure channel coincides with that of the one-logical-qubit zero-rate HaPPY code, the overall recovery performance is reduced for finite code instances: Though the asymptotic rate is zero for both codes, comparing instances at a fixed number of physical qubits leads to higher rate and lower logical recovery probability for the black-hole code.
@@
There are many ideas for future modification of LEGO_HQEC. Firstly, an extension for considering any hyperbolic tiling parameterized by $\{p,q\}$ would be enormously helpful, as it has been shown that certain classes of holographic codes have vastly different threshold behavior as the tiling is altered [steinberg2024far]. Furthermore, recent work [heterogeneous_holo_qrm, junyu_thesis] shows that incorporating multiple seed codes can allow for potentially universal fault-tolerant logical operations [heterogeneous_holo_qrm]; tools to support such fault tolerance studies could be very useful. One can also consider implementing functionality for greater control over the encoded logical qubits in the bulk, as recent work shows that recovery rates of finite-rate holographic codes greatly depend not only on the number $k$ of encoded logical qubits, but critically on the positions of said logical qubits in the bulk [alex_logical_recovery]. Even more broadly, it is known that the qLEGO formalism applies to non-Abelian stabilizer codes, such as the XP stabilizer variety [xp_stab_quantum_cao, webster2022xp], or even non-stabilizer codes, such as those presented in [exotic_gatesets, lidar2013quantum, non_additive1, non_additive2, non_additive3]; one may consider incorporating functionality to accommodate such holographic code constructions in the near future. Moreover, the class of codes known as Evenbly codes, a derivate of the hyperinvariant tensor networks first introduced in [hyper_evenbly, steinberg2022conformal], have been extensively studied using LEGO_HQEC [steinberg2024far, steinberg_2023]; however, one major impediment was the inclusion of Hadamard gates on the contracted edges of the tensor network, thereby hindering decoding of Pauli channels with the tensor-network decoder. Allowing for such constraints, among others, remains within the future goals of this work. Finally, there are many potential additions purely from a software and usability standpoint, such as: adding functionality for generalized concatenated codes [knill2005quantum, gottesman1997stabilizer, lidar2013quantum]; visualization tools, such as those that have recently appeared in the software package PlanqTN [pato_planqTN]; or GPU acceleration tools; among other prospects.
VI-A Acknowledgements
@@
We thank David Elkouss and Rob Harris for useful discussions. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation” and Berlin Quantum. CC acknowledges support from the Commonwealth Cyber Initiative at Virginia Tech.
VI-B Author Contributions